OpenStudy (anonymous):

What is the y-intercept of the line whose equation is y = 1/2 x - 3? -3 1/2 (faction) 3

OpenStudy (anteater):

Easy one for last! :D

OpenStudy (anteater):

y = mx + b :) Since b is your y-intercept ... what value does b have in the equation?

OpenStudy (anteater):

For example, if you had this linear equation: y = 3/4x + 7, m would be 3/4 and b would be 7

OpenStudy (anteater):

You have y = 1/2x - 3 ... so what is m and what is b for this equation?

OpenStudy (anonymous):

give me a second lol

OpenStudy (anteater):

k

OpenStudy (anteater):

Unfortunately, I must go. But for this equation, the slope m would be 1/2, and the y-intercept would be -3. I hope this was helpful!